The following is the balance sheet of Korver Supply Company at December 31, 2020 (prior year).

KORVER SUPPLY COMPANY

Balance Sheet

At December 31, 2020

Assets

Cash$160,000

Accounts receivable 320,000

Inventory 270,000

Furniture and fixtures (net) 180,000

Total assets$930,000

Liabilities and Shareholders’ Equity

Accounts payable (for merchandise)$270,000

Notes payable 280,000

Interest payable 7,000

Common stock 110,000

Retained earnings 263,000

Total liabilities and shareholders’ equity$930,000


Transactions during 2021 (current year) were as follows:

1.Sales to customers on account$930,000

2.Cash collected from customers 910,000

3.Purchase of merchandise on account 620,000

4.Cash payment to suppliers 630,000

5.Cost of merchandise sold 570,000

6.Cash paid for operating expenses 290,000

7.Cash paid for interest on notes 14,000


Additional Information:

The notes payable are dated June 30, 2020, and are due on June 30, 2022. Interest at 5% is payable annually on June 30. Depreciation on the furniture and fixtures for 2021 is $33,000. The furniture and fixtures originally cost $430,000.

Required:
Prepare a classified balance sheet at December 31, 2021, by updating ending balances from 2020 for transactions during 2021 and the additional information. The cost of furniture and fixtures and their accumulated depreciation are shown separately. (Amounts to be deducted should be indicated by a minus sign.)
Answer #1
Balance Sheet
Assets
Current Assets
Cash $    1,36,000 =160000+910000-630000-290000-14000
Accounts Receivable $    3,40,000 =320000+930000-910000
Inventory $    3,20,000 =270000+620000-570000
Total Current Assets $    7,96,000
Furniture and Fixture $    4,30,000
   Less : Accumulated Depreciation $   -2,83,000 =-(430000-180000)-33000
Total Land, Buildings and Equipment $    1,47,000
Total Assets $    9,43,000
Liabilities & Stockholder's Equity
Current Liabilities
Accounts Payable $    2,60,000 =270000+620000-630000
Interest Payable $          7,000 =7000-7000+7000
Total Current Liabilities $    2,67,000
Notes Payable $    2,80,000
Stockholder's Equity
Common Stock $    1,10,000
Retained Earnings $    2,86,000 =263000+23000
Total Stockholder's Equity $    3,96,000
Total Liabilities & Stockholder's Equity $    9,43,000
Income Statement
Revenue
Sales to customer $    9,30,000
Total Revenue $    9,30,000
Expenses
Cost of merchandise sold $    5,70,000
Operating Expenses $    2,90,000
Interest expense $       14,000
Depreciation expense $       33,000
Total Expenses $    9,07,000
Net Income $       23,000